OpenStudy (anonymous):

A seal went 15 feet below sea level to catch a fish. A sea lion dove 6 feet less than two times as deep as the seal to catch a larger fish. What expression represents the sea lion’s position in relation to sea level?

OpenStudy (anonymous):

x=2(15)-6 x=30-6 x=24

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Happy to help:)